OpenStudy (anonymous):

a = 40, c = 41, b = *DRAWING INC* 1 9 57 81

OpenStudy (anonymous):



OpenStudy (anonymous):

OOPS THE small b and c are the other way around.

OpenStudy (hoblos):

b= sqrt(41^2 - 40^2) =9

OpenStudy (anonymous):

what about a = 8, b = 15, c = same pic

OpenStudy (hoblos):

c= sqrt(8^2 +15^2) = 17

OpenStudy (anonymous):

thankyou=) ...1 more? a = 5, b = 12, c =

OpenStudy (hoblos):

can you try this one

OpenStudy (anonymous):

ok=) thanxx

OpenStudy (hoblos):

tell me what you get so i can check it for you

OpenStudy (anonymous):

154?

OpenStudy (hoblos):

not really... c=sqrt(5^2 + 12^2)=13

OpenStudy (anonymous):

hehe silly me, thats what i got on the calc. well thanxx..

OpenStudy (hoblos):

MY PLEASURE XD